# D Scaling Up Deepfake Detection by Learning from Discrepancy

### v1

@@ Abstract
The boom of Generative AI brings opportunities entangled with risks and concerns. In this work, we seek a step toward a universal deepfake detection system with better generalization and robustness, to accommodate the responsible deployment of diverse image generative models. We do so by first scaling up the existing detection task setup from the one-generator to multiple-generators in training, during which we disclose two challenges presented in prior methodological designs. Specifically, we reveal that the current methods tailored for training on one specific generator either struggle to learn comprehensive artifacts from multiple generators or tend to sacrifice their ability to identify fake images from seen generators (i.e., In-Domain performance) to exchange the generalization for unseen generators (i.e., Out-Of-Domain performance). To tackle the above challenges, we propose our Discrepancy Deepfake Detector (D) framework, whose core idea is to learn the universal artifacts from multiple generators by introducing a parallel network branch that takes a distorted image as extra discrepancy signal to supplement its original counterpart. Extensive scaled-up experiments on the merged UFD and GenImage datasets with six detection models demonstrate the effectiveness of our framework, achieving 5.3% accuracy improvement in the OOD testing compared to the current SOTA methods while maintaining the ID performance.
@@ 1 Introduction
Deepfake detection, which aims at discerning between real and generated fake images, acts as a crucial defense in responsibly deploying large generative models in the real world. The effectiveness of these detection systems largely depends on their generalization capability; ideally, they should reliably identify any synthesized image as fake, regardless of the generating source (i.e., the deep generative models from which they are generated). Accordingly, current deepfake detection methods [32; 45; 29; 5; 46; 10] usually follow a standard task setup for evaluating the performance, which involves training on images from one specific generator and testing on images from various unseen generators. As a natural step towards a universal deepfake detection foundation model, in this work, we propose to scale up this current setup from “train-on-one and test-on-many” to “train-on-many and test-on-many”, to accommodate the rapid development pace from the generative modeling side and better align with the real-world scenario.
@@
However, in the process of scaling up existing deepfake detection models to multiple generators, we encounter two challenges that have not yet been adequately addressed in prior literature, limiting the generalization performance and robustness of an expanded detection system. First, we reveal that the existing methodology designs tailored for the “train-on-one and test-on-many” setup struggle to learn the comprehensive and universal artifacts when the fake images in training present different fingerprint patterns from more than one generator. For instance, one of the current SOTA detection models, DIRE from ICCV’23 [46], tends to learn the most obvious generator-specific artifacts (e.g., fingerprints of diffusion models [21; 42]) but overlooks the subtle but invariant common artifacts shared by other generative AI models (e.g., fingerprints from GAN-based generators [17]), leading to inferior testing performance on the images produced by unseen generators. As detailed in Sec. 4, our experimental results show that, when scaling up to 8 training generators and 12 unseen generators, DIRE [46] achieves 97.6% on in-domain (ID) testing, but performs significantly worse on the critical out-of-domain (OOD) testing with only 68.4% accuracy. Second, our empirical analysis further suggests that prior methods tend to comprise the trade-off between training and testing performance, by “underfitting” the ID training to barter for better OOD generalization ability in testing. As a concrete example, another SOTA detection method, UFD from CVPR’23 [32], achieves better generalization on OOD testing with an accuracy of 81.4% compared to DIRE [46]. However, the better OOD performance of UFD comes with the expense of the lower ID testing performance, resulting in only 86.6% on the ID accuracy.
@@
To tackle the above challenges on the way of scaling up the deepfake detection methods to a universal system, we introduce our D framework, terminology from Discrepancy Deepfake Detector, for an extended setup under the “train-on-many and test-on-many” scenario. The secret recipe of D lies within the core idea of exploiting and learning the universal artifacts among various deep generators, which intuitively facilitates the learning and improves the testing robustness. Specifically, unlike the existing methods [45; 5; 32] that feed the detection model only with either real or generated images, we adopt a two-branch design to provide D with an extra corrupted image corresponding to its original counterpart, as illustrated in Fig. 1 and detailed in Sec. 3. The distorted image, which is obtained via operations such as patch shuffling, flipping, and rotation, deconstructs the unique fingerprint from a specific generator and serves as a discrepancy signal to supplement the original image. This unique methodological design encourages the proposed D framework to effectively capture the invariant artifacts from the image pair. We further improve the robustness and generalization ability of our D by extracting the image features from the CLIP encoder [34], and enhance their interactions via self-attention layers before making the final prediction.
@@
We perform extensive experiments by gradually scaling up six deepfake detection methods (including [32; 5; 29; 45; 46] and our proposed D) from conventional setup to 8 training generators and 12 testing unseen generators, on the UniversalFakeDetect (UFD) [32] and the GenImage [52] datasets. Each detection model is trained till convergence to ensure a fair comparison. Our comprehensive experiments and analysis not only explicitly disclose the two previously mentioned issues, but also demonstrate the effectiveness of our proposed D by outperforming SOTA methods with 5.3% on the OOD testing mean accuracy.
@@
We propose to scale up the current deepfake detection setup from one-generator to multiple-generators in training to accommodate the fast development pace from the generative modeling side and disclose two factors in the existing methods that limit the generalization ability and robustness in this extended detection scenario.
We introduce our Discrepancy Deepfake Detector (D) framework to address the above challenges, whose core idea is to learn the general artifacts among multiple generators with an extra distorted image as the discrepancy signal via a parallel network branch.
We perform extensive experiments, by gradually scaling up the current setup to 8 generators, and achieve 5.3% improvement in the OOD testing performance compared to SOTA methods, while maintaining the ID performance.
@@ 2 Related Work
With the swift advancement of generative models [18; 27; 21; 43; 36; 13; 39] and customization methods [38; 16; 47; 53; 26; 49], synthesized images are becoming increasingly realistic, making it challenging for the human eye to differentiate between real and fake images. To counter potential malicious usage, a considerable amount of research has been devoted to detecting generated images, yielding impressive identification accuracy.
@@ 2.1 Artifact-based Detection Methods
These methods explicitly extract artifacts in generated images in a preprocessing way. They usually directly estimate the artifacts [50; 22], compute the residual noise between the original image and its reconstructed one [30; 29; 46], or calculate gradients [44]. [30] was the first to reveal that GANs leave specific fingerprints on the generated images. This was achieved by extracting the noise residual through an appropriate denoising filter, demonstrating its potential for forensic analysis. [50] proposed a GAN simulator to emulate the artifacts produced by the common pipeline shared by several popular GAN models. [22] proposed a fingerprint generator to cover the fingerprints of various GAN models to achieve generalized detection ability and avoid data dependency. LNP [29] pretrain a reconstruction network with only real images to extract the Learned Noise Patterns (LNP), which can explicitly embody artifacts. DIRE [46] used the prior from a pre-trained diffusion model to determine whether an image is within the distribution of fake images. This was achieved by computing the Diffusion Reconstruction Error (DIRE). [44] employed a pre-trained CNN model to extract the gradients of target images as the representation of artifacts.
@@ 2.2 Learning-based Detection Methods
This type of method considers directly learning the artifacts from raw images. They aim to design a better network to extract more discriminative features. CNNDet [45] directly trained ResNet50 as the classifier on their proposed semantic-aligned dataset, containing fake images generated by ProGAN. Their methods exhibited strong generalization capabilities to other CNN-based generative models. Patchfor [5] used a fully convolutional patch-based classifier, emphasizing local patches over global structure. [19] modified the ResNet50 network with two fewer down-sampling layers to improve the detection performance. However, the learning-based methods directly train the entire network to learn the presence or absence of actual artifact patterns. They often overlook features inherent to real images, which can lead to overfitting and poor generalization.
2.3 CLIP-based Detection Methods
To mitigate the shortcomings of the learning-based methods, this type of method leverages the frozen feature space of CLIP to extract more generalizable features. UFD [32] first highlighted that a feature space not explicitly learned for generated image detection is more effective due to its unbiased decision boundaries. Consequently, they performed detection in the pre-trained CLIP feature space using linear probing or nearest neighbor methods, demonstrating impressive generalization performance on unseen generated images. Taking a further step in this direction, [10] trained SVMs with the features from the penultimate layer and achieved commendable performance despite a limited amount of well-organized training data. Likewise, [41] additionally utilized the text encoder of the pre-trained CLIP model to exploit the prompt information by concatenating the encoded image embeddings and text embeddings together as features. However, these CLIP-based methods employ only one linear classifier to distinguish between real and fake features of images. Therefore, when applied to a larger multi-generator dataset, the simple classifier struggles to effectively fit the ID data, which further impacts OOD performance. In this work, we further enhance the fitting and generalization ability of CLIP-based detectors on multi-generator datasets. We propose to learn more stable and invariant artifacts by comparing images before and after corruption. Our approach facilitates a better understanding of the universal differences between real and fake images.
@@ 3.1 Problem Formulation
Formally, given a generator set composed of different types of generators, , and denote the real and fake samples of , where is the number of samples in each class. Our goal is to correctly distinguish between real and fake when given any image from the overall dataset, . The existing detection task setup involves training on one of the generators . In terms of evaluation, they regard the seen generator as the ID generator and the other unseen generators as the OOD generators. This setup implies learning to detect universal artifacts from only one specific kind of artifact, which is hard and unnecessary. This impractical setup may mislead the research community because the generalization performance is unstable and would highly depend on the choice of the training generator. For example, the OOD performances of UFD [32] would drop significantly from 87.6% to 58.9% when switching the training from the ProGAN [23] to Midjourney [2] in this scheme.
Therefore, we propose to step out of the “train-on-one and test-on-many” task setup and leverage a new “train-on-many and test-on-many” evaluation scheme. More specifically, we add more generators into the current dataset and train on a subset composed of generators, denoted by . In this way, we regard the generators in as the ID generators and the other generators as the OOD generators. In practice, we merge two large-scale existing deepfake detection datasets, i.e., the UFD dataset [32] and GenImage dataset [52]. Finally, there are in total 20 state-of-the-art generators available in the scale-up setup. Considering the diversity, we regard all the 8 generators used in different previous works, including 2 GANs methods like ProGAN [23], and 6 DMs like ADM [13]. The rest 12 generators (including both GANs and DMs) are used as the OOD set.
@@ 3.2 Challenges in the Scaled-up Setup
Based on the scaled-up setup, we gradually add the generators into the training pool in different random orders to evaluate the existing works. As the number of generators increased from one to eight, we found existing methodology designs tailored for the previous setup could not learn universal artifacts. For example, one of the state-of-the-art works UFD [32] from CVPR’23 even decreases significantly in terms of ID testing accuracy as more generators are used for training (Fig. 2), demonstrating the underfitting issue. While another state-of-the-art work DIRE [46] from ICCV’23 shows extremely good ID performance (97.6%), it suffers from the generalization issue that the OOD testing accuracy is not promising (68.4%), despite it has already seen eight advanced generators covering typical GANs and DMs. These results suggest existing models could not find the invariant artifacts that are universal in multiple generators. The main reason might be the universal data patterns we wish the model to learn are coupled with or hidden by semantics. It inspires us to go deeper to learn universal artifacts rather than the superficial and generator-specific artifacts that can be directly seen on the deepfake images.
3.3 D: Deepfake Dectection by Learning from Discrepancy
@@
Visual Backbone. Directly learning on the deepfake detection data without pretraining can lead to overfitting on the training generators, resulting in poor generalization to OOD generators [9; 52]. To mitigate this, we follow previous work [32; 10] and adopt the pre-trained CLIP:ViT-L/14 as our feature extractor. More specifically, we extract features from the penultimate layer of the pre-trained CLIP model, which keeps more detailed visual clues of the input.
@@
where and are the visual embeddings of the respective images, denotes the extraction of the penultimate feature using the pre-trained CLIP model, PatchShuffle is the patch-shuffling operation, and AUG represents the common data augmentation method used in the previous work [45].
Self-Attention Invariance Extraction. Having obtained two features, and , from the original and patch-shuffled images respectively, we stack the two features together, . The correlation between and is crucial for detecting fake images. To this end, we employ a self-attention layer on the stacked feature to extract the shared artifact feature by learning the associations between the features. In this way, we force the model to learn from the discrepancy between the original image and its patch-shuffled image. In experiments, we found the more distinct between the original image and the corrupted image, the better the deepfake detection performance the model can get.
@@
where is the batch size, and is the label of the -th input sample. Note that our visual backbone is frozen, thus the learnable parameters are only self-attention and the classification layer.

### v2

# D3: Scaling Up Deepfake Detection by Learning from Discrepancy

@@ Abstract
The boom of Generative AI brings opportunities entangled with risks and concerns. Existing literature emphasizes the generalization capability of deepfake detection on unseen generators, significantly promoting the detector’s ability to identify more universal artifacts. This work seeks a step toward a universal deepfake detection system with better generalization and robustness. We do so by first scaling up the existing detection task setup from the one-generator to multiple-generators in training, during which we disclose two challenges presented in prior methodological designs and demonstrate the divergence of detectors’ performance. Specifically, we reveal that the current methods tailored for training on one specific generator either struggle to learn comprehensive artifacts from multiple generators or sacrifice their fitting ability for seen generators (i.e., In-Domain (ID) performance) to exchange the generalization for unseen generators (i.e., Out-Of-Domain (OOD) performance). To tackle the above challenges, we propose our Discrepancy Deepfake Detector (D) framework, whose core idea is to deconstruct the universal artifacts from multiple generators by introducing a parallel network branch that takes a distorted image feature as an extra discrepancy signal and supplement its original counterpart. Extensive scaled-up experiments demonstrate the effectiveness of D, achieving 5.3% accuracy improvement in the OOD testing compared to the current SOTA methods while maintaining the ID performance. The source code will be updated in our GitHub repository: https://github.com/BigAandSmallq/D3.
@@ 1 Introduction
Deepfake detection, which aims at discerning between real and generated fake images, acts as a crucial defense in responsibly deploying large generative models in the real world. The effectiveness of these detection systems largely depends on their generalization capability; ideally, they should reliably identify any synthesized image as fake, regardless of the generating source (i.e., the deep generative models from which they are generated). Accordingly, current deepfake detection methods [30, 42, 27, 5, 43, 10] usually follow a standard task setup for evaluating performance, which involves training on images from one specific generator and testing on images from various unseen generators, emphasizing and enhancing the detector’s ability to identify universal artifacts from a single generator source. As a natural step towards a universal deepfake detection foundation model, in this work, we propose to scale up this current setup from “train-on-one and test-on-many” to “train-on-many and test-on-many”, to accommodate the rapid development pace from the generative modeling side and better align with the real-world scenario.
@@
However, in the process of scaling up existing deepfake detection models to multiple generators, existing methods encounter two challenges that have not yet been adequately addressed in prior literature, limiting the generalization performance and robustness of an expanded detection system. First, we reveal that the existing methodology designs tailored for the “train-on-one and test-on-many” setup struggle to learn the comprehensive and universal artifacts when the fake images in training present different fingerprint patterns from more than one generator. For instance, one of the current SOTA detection models, DIRE from ICCV’23 [43], tends to learn the most obvious generator-specific artifacts (e.g., fingerprints of diffusion models) but overlooks the subtle but invariant common artifacts shared by other generative AI models (e.g., fingerprints from GAN-based generators), leading to inferior testing performance on the images produced by unseen generators. As detailed in Sec. 4, our experimental results show that, when scaling up to 8 training generators and 12 unseen generators, DIRE [43] achieves 97.6% on in-domain (ID) testing, but performs significantly worse on the critical out-of-domain (OOD) testing with only 68.4% accuracy. Second, our empirical analysis further suggests that prior methods tend to comprise the trade-off between training and testing performance, by “underfitting” the ID training to barter for better OOD generalization ability in testing. As a concrete example, another SOTA detection method, UFD from CVPR’23 [30], achieves better generalization on OOD testing with an accuracy of 81.4% compared to DIRE [43]. However, the better OOD performance of UFD comes with the expense of the lower ID testing performance, resulting in only 86.6% on the ID accuracy.
Furthermore, Detectors exhibit generalization performance divergence when scaled from one to eight generators. As illustrated in Fig. 2, Patchfor [5], initially slightly weaker than CNNDet [42], greatly surpasses CNNDet by 9% after scaling up. This suggests that the proposed new setup intrinsically demands a comprehensive ability to acquire universal image artifacts across multiple generators for achieving higher generalization, which is different compared to the current setup.
@@
To tackle the above challenges on the way of scaling up the deepfake detection methods to a universal system, we introduce our D framework, terminology from Discrepancy Deepfake Detector, for an extended setup under the “train-on-many and test-on-many” scenario. The secret recipe of D lies within the core idea of deconstructing the universal artifacts from the manually introduced discrepancy in image features, which intuitively facilitates the universal artifacts learning among various generators and improves the testing robustness. Specifically, unlike the existing methods that feed the detection model only with either real or generated images, we adopt a two-branch design to provide D with an extra distorted image corresponding to its original counterpart, as illustrated in Fig. 1 and detailed in Sec. 3.
Intuitively, the distorted image, created via operations like patch shuffling, flipping, and rotation, deconstructs the unique fingerprint from a specific generator into a different level of view and serves as a discrepancy signal to supplement the original image to find the generic artifact across generators. This unique methodological design encourages the proposed D framework to capture the different artifacts distributed in the original and distorted images respectively, and the artifact correlation between them.
@@
We perform extensive experiments by gradually scaling up six deepfake detection methods (including [30, 5, 27, 42, 43] and our proposed D) from conventional setup to 8 training generators and 12 testing unseen generators, on the UniversalFakeDetect (UFD) [30] and the GenImage [50] datasets. Each detection model is re-trained till convergence to ensure a fair comparison. Our comprehensive experiments and analysis explicitly disclose the two previously mentioned issues by outperforming SOTA methods with 5.3% on the OOD testing mean accuracy while maintaining the ID performance.
@@
We propose to scale up the current deepfake detection setup from one-generator to multiple-generators in training to accommodate the real-world scenario and disclose two factors in the existing methods that limit the generalization ability and robustness in this extended detection scenario.
We perform extensive experiments, by gradually scaling up 5 baselines in the current setup to 8 generators, showing the intrinsic requirements for achieving higher generalization ability in this new setup are different from the current setup.
We introduce a novel framework Deepfake Detector (D) to address the above challenges, whose core idea is to learn the general artifacts among multiple generators with an extra distorted image as the discrepancy signal via a parallel network branch. Our method achieves 5.3% improvement in the OOD testing performance compared to SOTA methods while maintaining the ID performance.
@@ 2 Related Work
With the swift advancement of generative models [17, 25, 19, 40, 34, 13, 37] and customization methods [36, 16, 44, 51, 24, 47], synthesized images are becoming increasingly realistic, making it challenging for the human eye to differentiate between real and fake images. To counter potential malicious usage, a considerable amount of research has been devoted to detecting generated images, yielding impressive identification accuracy.
@@ 2.1 Artifact-based Detection Methods
These methods explicitly extract artifacts in generated images in a preprocessing way. They usually directly estimate the artifacts [48, 20], compute the residual noise between the original image and its reconstructed one [28, 27, 43], or calculate gradients [41]. [28] was the first to reveal that GANs leave specific fingerprints on the generated images. This was achieved by extracting the noise residual through an appropriate denoising filter, demonstrating its potential for forensic analysis. LNP [27] pretrain a reconstruction network with only real images to extract the Learned Noise Patterns (LNP), which can explicitly embody artifacts. DIRE [43] used the prior from a pre-trained diffusion model to determine whether an image is within the distribution of fake images. This was achieved by computing the Diffusion Reconstruction Error (DIRE).
@@ 2.2 Learning-based Detection Methods
This type of method considers directly learning the artifacts from raw images. They aim to design a better network to extract more discriminative features. CNNDet [42] directly trained ResNet50 as the classifier on their proposed semantic-aligned dataset, containing fake images generated by ProGAN. Their methods exhibited strong generalization capabilities to other CNN-based generative models. Patchfor [5] used a fully convolutional patch-based classifier, emphasizing local patches over global structure. However, the learning-based methods directly train the entire network to learn the presence or absence of actual artifact patterns. They often overlook features inherent to real images, which can lead to overfitting and poor generalization.
2.3 Pretrained-based Detection Methods
To mitigate the shortcomings of the learning-based methods, this type of method leverages the frozen feature space to extract more generalizable features. UFD [30] first highlighted that a feature space not explicitly learned for generated image detection is more effective due to its unbiased decision boundaries. Consequently, they performed detection in the pre-trained CLIP feature space using linear probing or nearest neighbor methods, demonstrating impressive generalization performance on unseen generated images. Taking a further step in this direction, [10] trained SVMs with the features from the penultimate layer and achieved commendable performance despite a limited amount of well-organized training data. Likewise, [39] additionally utilized the text encoder to exploit the prompt information by concatenating the encoded image embeddings and text embeddings together as features. However, these current Pretrained-based methods heavily rely on linear separability in the feature space to distinguish between real and fake images. Therefore, when applied to a larger multi-generator dataset, the single pre-trained feature space and simple classifier struggle to effectively fit the more complex ID data, which further impacts OOD performance. In this work, we further enhance the fitting and generalization ability of Pretrained-based detectors on multi-generator datasets. We achieve this by supplementing a distorted image signal to deconstruct the artifacts and learn more universal and invariant artifacts, facilitating a better understanding of the universal artifacts.
@@ 3.1 Problem Formulation
Formally, given a generator set composed of different types of generators, , and denote the real and fake samples of , where is the number of samples in each class. Our goal is to correctly distinguish between real and fake when given any image from the overall dataset, . The existing detection task setup involves training on one of the generators . In terms of evaluation, they regard the seen generator as the ID generator and the other unseen generators as the OOD generators. This setup implies learning to detect universal artifacts from only one specific kind of artifact, which is hard and unnecessary. This impractical setup may mislead the research community because the generalization performance is unstable and would highly depend on the choice of the training generator. For example, the OOD performances of UFD [30] would drop significantly from 83.2% to 58.9% when switching the training from the ADM [13] to Midjourney [2] in this scheme.
Therefore, we propose to step out of the “train-on-one and test-on-many” task setup and leverage a new “train-on-many and test-on-many” evaluation scheme. Specifically, we add more generators into the current dataset and train on a subset composed of generators, denoted by . In this way, we regard the generators in as the ID generators and the other generators as the OOD generators. In practice, we merge two large-scale existing deepfake detection datasets, i.e., the UFD dataset [30] and GenImage dataset [50]. Finally, there are in total 20 state-of-the-art generators available in the scale-up setup. Considering the diversity, we regard all the 8 generators used in different previous works, including 2 GANs methods like ProGAN [21], and 6 DMs like ADM [13]. The rest 12 generators (including both GANs and DMs) are used as the OOD set.
@@ 3.2 Challenges in the Scaled-up Setup
Based on the scaled-up setup, we gradually add the generators into the training pool in different random orders to evaluate the existing works. As the number of generators increased from one to eight, we found existing methodology designs tailored for the previous setup could not learn universal artifacts. For example, one of the state-of-the-art works UFD [30] from CVPR’23 even decreases significantly in terms of ID testing accuracy as more generators are used for training (Fig. 2), demonstrating the underfitting issue. While another state-of-the-art work DIRE [43] from ICCV’23 shows extremely good ID performance (97.6%), it suffers from the generalization issue that the OOD testing accuracy is not promising (68.4%), despite it has already seen eight advanced generators covering typical GANs and DMs. These results suggest existing models could not find the invariant artifacts that are universal in multiple generators. The main reason might be these detectors struggle to capture the universal multi-generator artifact patterns we wish the model to learn.
In addition, detectors exhibit performance divergence when scaled from one to eight generators. For example, Patchfor, initially weaker than CNNDet and just comparable to LNP and DIRE, achieves a substantial increase to 0.789 when using eight generators. This performance greatly surpasses CNNDet (0.699), LNP (0.719), and DIRE (0.684). It indicates that the intrinsic requirements for achieving higher generalization under multi-generator training are different from the one-generator training. The scaled-up setup requires a more comprehensive ability to acquire universal image artifacts across multiple generators. This inspires us to go deeper to deconstruct the artifacts of each specific generator into different levels of view, therefore learning the universal artifacts rather than the superficial and generator-specific artifacts that can be directly seen on the deepfake images.
3.3 D: Discrepancy-based Deepfake Detection
@@
Visual Backbone. Directly learning on the deepfake detection data without pretraining can lead to overfitting on the training generators, resulting in poor generalization to OOD generators [9, 50]. To mitigate this, we follow previous work [30, 10] and adopt the pre-trained CLIP:ViT-L/14 as our feature extractor. More specifically, we extract features from the penultimate layer of the pre-trained CLIP model, which keeps more detailed visual clues of the input.
@@
where and are the visual embeddings of the respective images, denotes the extraction of the penultimate feature using the pre-trained CLIP model, PS is the patch-shuffling operation, and AUG represents the common data augmentation method used in the previous work [42].
Self-Attention Invariance Extraction. Having obtained two features, and , from the original and patch-shuffled images respectively, we stack the two features together, . The correlation between and is crucial for detecting fake images. To this end, we employ a self-attention layer on the stacked feature to extract the shared artifact feature by learning the associations between the features. In this way, we force the model to learn from the discrepancy between the original image and its patch-shuffled image. In experiments, we find that the more distinct the original image and the corrupted image are, the better the deepfake detection performance of the model is.
@@
where is the batch size, and is the label of the -th input sample. Our backbone is frozen, thus the learnable parameters are only self-attention and the classification layer.